Line Integral & Path Independency Problem 1 Prove that the vector field F = (2x – 3yz?) { +(2 – 3xz) j-6xyzk is the gradientFind the function whose gradient is F. For these two vectors Ħf and to be equal, the first, second, and third terms in one veThis is all one question please answer asap
Answer #1

Answers F = (2x - 34 22)i + (2-3x22) ĵ - (Gry z) Å L Curl = NxF Z d d dx sy d2 - 6x42 by Curl F SY 2x-3y22 2-3x22 Solving thef = J Fy dy 12- 2- 3 x 22) dy at f - 2Y - 34x22 + df đz NY - - 6xyz - f = (-bayz dz f 3 xyz². * Thus, by grouping without rep